When probabilities are written in terms of x, use the fact that they add to 1, or the given total probability, to form an equation in x, then solve it.
Write each probability in terms of x.
Form an equation: all outcomes add to 1, or a product along branches equals the given value.
Solve, and reject any value that gives a probability outside 0 to 1.
Watch out Multiply along a branch using each probability given what came before it.
See it worked
Worked exampleNon-calculator·3 marks
An experiment has exactly four possible outcomes A, B, C and D. Their probabilities are 2x, 3x, x + 0.1 and 0.3 respectively. Work out the probability of outcome B.
2x+3x+x+0.1+0.3=1
6x=0.6
The probabilities add to 1: 2x + 3x + x + 0.1 + 0.3 = 1.
Therefore 6x = 0.6, so x = 0.1.
P(B) = 3x = 0.3.
Answer 0.3
Your turn
Question 1Non-calculator·3 marks
P(A)=0.4,P(B)=0.5 and P(A∩B)=0.2.
(a) Are A and B independent? You must show how you get your answer.
(b) Work out P(A∪B).
Question 2Non-calculator·5 marks
A bag contains x red counters and 5 blue counters, and there are more blue counters than red. Two counters are taken at random without replacement. The probability that they are different colours is 95.
Work out the value of x.
Question 3Non-calculator·3 marks
A bag contains n red counters and 4 blue counters, where n is greater than 4. Two counters are selected at random without replacement. The probability that the counters are different colours is 8/15. Work out n. You must show your working.
In past papers
16 official questions in 16 papers since 2019. The latest: